Witt kernels of function field extensions in characteristic 2

Ahmed Laghribi^{1}

Université d’Artois Faculté des Sciences Jean Perrin Rue Jean Souvraz SP-18 F-62307 Lens, France

Abstract

In characteristic 2 we give a complete characterization of anisotropic symmetric bilinear forms that become metabolic over the function field of a quadratic form. We also study the hyperbolicity of nonsingular quadratic forms over such a field by generalizing some results by Fitzgerald [6]. As an application, we introduce and study the notion of Pfister neighbors for bilinear forms, and classify anisotropic bilinear forms of height 1, i.e. those that become metabolic over their own function fields. Other consequences are also included.

Key words: Bilinear and quadratic forms, hyperbolicity, metabolicity, Witt group and Witt ring of a field, Witt kernels of field extensions. 2000 MSC: Primary 11E04, Secondary 11E81

1 Introduction

Let F denote a field of characteristic 2, and let W_{q}(F) (resp. W (F)) be the Witt group of nonsingular quadratic forms over F (resp. the Witt ring of regular symmetric bilinear forms over F). For any field extension K/F, there are homomorphisms W_{q}(F) ⟶^{i} W_{q}(K) and W (F) ⟶^{j} W (K) induced by scalar extension.

  Email address: laghribi@euler.univ-artois.fr (Ahmed Laghribi).

1 The author was supported by the European research network HPRN-CT-2002-00287 “Algebraic K-Theory, Linear Algebraic Groups and Related Structures”.

An important problem in the algebraic theory of quadratic and bilinear forms is the study of the kernels W_{q}(K/F) and W (K/F) of the homomorphisms i and j, respectively.

In this paper we are interested to the case where K = F (φ), the function field of a quadratic form φ. For such a field in characteristic different from 2, the study was done by Fitzgerald [6]. Our aim is to treat the case of characteristic 2. More precisely, we will give a complete computation of W (F (φ)/F), and for the kernel W_{q}(F (φ)/F) we will extend important results in [6]. Our study includes function fields of possibly singular quadratic forms.

1.1 The kernel W (F (φ)/F)

Throughout this paper, the expression “bilinear form” means “regular symmetric bilinear form”.

For a quadratic form (resp. a bilinear form) φ with underlying vector space V, we denote by D_{F} (φ) the set of scalars in F^{∗} represented by φ (resp. the set of scalars φ(v, v) ∈ F^{∗} for v ∈ V). A totally singular form of dimension n is a quadratic form isometric to ∑_{i=1}^{n} a_{i}x_{i}^{2} for some a_{i} ∈ F^{∗}. We denote it by ⟨a_{1}, ⋯, a_{n}⟩. A bilinear form B is called associated to a totally singular form φ (or φ is associated to B) if dim B = dim φ and D_{F} (B) = D_{F} (φ). Note that D_{F} (B) is nothing but the F^{2}-vector space inside F generated by the elements in the diagonal of any matrix representing B, and that dim_{F^{2}} D_{F} (B) ≤ dim B with equality if B is anisotropic. Moreover, two totally singular forms associated to the same bilinear form are isometric. We denote by B̃ the totally singular form associated to a bilinear form B, and by A(φ) the set of bilinear forms associated to a totally singular form φ.

Let ⟨a_{1} : b : a_{2}⟩ denote the 2-dimensional bilinear form whose underlying vector space has a basis {e_{1}, e_{2}} satisfying: B(e_{i}, e_{i}) = a_{i} and B(e_{1}, e_{2}) = b. A metabolic plane is a 2-dimensional bilinear form isomorphic to ⟨a : 1 : 0⟩ for some a ∈ F. A metabolic bilinear form is an orthogonal sum of metabolic planes. Recall that a bilinear form is isotropic if and only if it contains a metabolic plane as a subform.

Our first result in this paper is the following proposition which gives the subform theorem version for bilinear forms. It will be very useful to study the metabolicity of bilinear forms over function fields of quadratic forms:

Proposition 1.1 Let B be an anisotropic bilinear form, and let φ be an anisotropic quadratic form such that B is metabolic over F (φ). Then φ is totally singular and for any α ∈ D_{F} (φ)D_{F} (B), there exists a subform B′ of αB such that B′ ∈ A(φ). In particular, dim φ ≤ dim B.

We also need the notion of norm field introduced in [7]. Recall that the norm field of a nonzero totally singular form φ, denoted by N_{F} (φ), is the field F^{2}(ab | a, b ∈ D_{F} (φ)). The degree of the extension N_{F} (φ)/F^{2}, denoted by ndeg_{F} (φ), is called the norm degree of φ. If b_{1}, ⋯, b_{m} ∈ F^{∗} are such that N_{F} (φ) = F^{2}(b_{1}, ⋯, b_{m}) and ndeg_{F} (φ) = 2^{m}, then we denote by φ_{qp} the totally singular form associated to the anisotropic m-fold bilinear Pfister form ⟨1, b_{1}⟩_{b}⊗⋯⊗⟨1, b_{m}⟩_{b}, where ⟨a_{1}, ⋯, a_{n}⟩_{b} denotes the bilinear form ∑_{i=1}^{n} a_{i}x_{i}y_{i} for a_{i} ∈ F^{∗}. We prove that the anisotropic part φ_{an} of φ is similar to a subform of φ_{qp}. Moreover, if φ_{an} is similar to a subform of a totally singular form φ′ which is associated a bilinear Pfister form, then φ_{qp} is a subform of φ′. For more details on norm fields we refer to [7, Section 8].

Now our criterion concerning the metabolicity of bilinear forms over function fields of quadratic forms is as follows:

Theorem 1.2 Let φ be an anisotropic quadratic form of dimension ≥ 2 such that W (F (φ)/F) ≠ 0. Then φ is totally singular, and an anisotropic bilinear form B over F becomes metabolic over F (φ) if and only if B ≃ α_{1}B_{1} ⊥ ⋯ ⊥ α_{r}B_{r} for some α_{1}, ⋯, α_{r} ∈ F^{∗} and B_{1}, ⋯, B_{r} d-fold bilinear Pfister forms belonging to A(φ_{qp}), where d is such that ndeg_{F} (φ) = 2^{d} (≃ and ⊥ denote isometry and orthogonal sum, respectively).

1.2 The kernel W_{q}(F (φ)/F)

Contrary to the case of bilinear forms, the hyperbolicity of nonsingular quadratic forms was previously studied over some field extensions, like quadratic and biquadratic extensions. This was done by Baeza [3], [4, Cor. 4.16, Page 128], Mammone and Moresi [17], and Hamza [1, Cor. 2.8], [2]. Recently, the author gave a complete computation of W_{q}(K/F) for K/F purely inseparable multiquadratic extension [16].

As was investigated in characteristic different from 2 by Elman, Lam, Wadsworth and Fitzgerald [5], [6], an important question related to the computation of W_{q}(F (φ)/F) is to know if these kernels are generated by Pfister forms (up to scalars), and whether the description can be given up to isometry instead of Witt-equivalence. In this sense, we recall the terminology of “Pfister group” and “strong Pfister group” originally due to Elman, Lam and Wadsworth:

Definition 1.3 For an integer n ≥ 1, we denote by P_{n}F the set of quadratic forms isometric to n-fold Pfister forms, and GP_{n}F = F^{∗}P_{n}F. For a field extension K/F and a subset I of ℕ, we say that:

(1) W_{q}(K/F) is an I-Pfister group if it is generated by quadratic forms in W_{q}(K/F) ∩ GP_{n}F for n ∈ I.

(2) W_{q}(K/F) is a strong I-Pfister group if any quadratic form in W_{q}(K/F) is isometric to an orthogonal sum of quadratic forms in W_{q}(K/F)∩GP_{n}F for n ∈ I.

The following theorem describes, up to isometry, nonsingular quadratic forms that become hyperbolic over the function field of a Pfister neighbor or quasi-Pfister neighbor (cf. subsection 2.3). In particular, for the quasi-Pfister neighbors case, this generalizes the result [1, Cor. 2.8] by Hamza:

Theorem 1.4 Let φ be an anisotropic Pfister neighbor or quasi-Pfister neighbor of π. Then we have the following:

(1) W_{q}(F (φ)/F) = W_{q}(F (π)/F).

(2) If φ is a Pfister neighbor, then any anisotropic form in W_{q}(F (φ)/F) is isometric to τ ⊗ π for some bilinear form τ. In particular, W_{q}(F (φ)/F) is a strong n-Pfister group, where n satisfies dim π = 2^{n}.

(3) If φ is a quasi-Pfister neighbor, then any anisotropic form in W_{q}(F (φ)/F) is isometric to B ⊗ ρ for some nonsingular quadratic form ρ, where B is a bilinear Pfister form satisfying B̃ ≃ π. In particular, W_{q}(F (φ)/F) is a strong (n + 1)-Pfister group, where n is as in (2).

However, in general, we have no criterion to compute W_{q}(F (φ)/F) for an arbitrary quadratic form φ. The principal idea that we will use to get informations on the kernel W_{q}(F (φ)/F) is to compare it to another one W_{q}(F (φ′)/F) provided that φ′ satisfies some properties. Our main result in this sense is Theorem 1.5 which treats the case when φ′ is dominated by φ, dim φ = dim φ′ + 1 and W_{q}(F (φ′)/F) is a strong n-Pfister group for some n ≥ 1 (dominated means that φ′ is the restriction of φ to a subspace of its underlying vector space; cf. subsection 2.2):

Theorem 1.5 Let φ be an anisotropic quadratic form (possibly singular) of dimension ≥ 3, and let φ′ be a quadratic form dominated by φ with codimension 1, i.e. dim φ = dim φ′ + 1, such that W_{q}(F (φ′)/F) is a strong n-Pfister group for some n ≥ 1. Then W_{q}(F (φ)/F) is an {n, n+1}-Pfister group. Moreover, if W_{q}(F (φ)/F) ∩ P_{n}F = {0}, then W_{q}(F (φ)/F) is an (n + 1)-Pfister group.

Theorem 1.5 partially extends [6, Prop. 1.2] to characteristic 2, and can be refined in the case of a Pfister neighbor φ′ as follows:

Theorem 1.6 Let φ be an anisotropic quadratic form of dimension ≥ 3, and let φ′ be a quadratic form dominated by φ with codimension 1. If φ′ is a Pfister neighbor of an n-fold Pfister form, then W_{q}(F (φ)/F) is a strong m-Pfister group where m = n or n + 1 according as φ is a Pfister neighbor of an n-fold Pfister form or not.

This paper is organized as follows. In the next section, we recall definitions and notions on bilinear and quadratic forms. Sections 3 is devoted to some preliminaries that we need for the proofs of our results, and then in section 4 we give our proofs. In view of the subform theorem stated in Proposition 1.1, we introduce in section 5 the notion of Pfister neighbors for bilinear forms. We prove preliminary results on such forms, similar to those known for Pfister neighbors in the case of quadratic forms. By using Theorem 1.2, we prove that anisotropic bilinear forms of height 1 are those similar to bilinear Pfister forms (Corollary 5.5). We also give a characterization of Pfister neighbors based on a splitting property over their own function fields, by proving that an anisotropic bilinear form B is a Pfister neighbor if and only if there exists B′ ∈ A(B̃) such that the anisotropic part of B′_{F (B)} is defined over F (Corollary 5.6). This is similar to a result by Knebusch for Pfister neighbors in characteristic different from 2 [12], and its generalization to characteristic 2 by the author and Hoffmann [7]. In the last section, we give further results on Witt kernels, and we give a detailed description of W_{q}(F (φ)/F) for φ of small dimension.

2 Backgrounds on bilinear and quadratic forms

Any quadratic form φ of dimension ≥ 1 can be written up to isometry

φ ≃ [a_{1}, b_{1}] ⊥ ⋯ ⊥ [a_{r}, b_{r}] ⊥ ⟨c_{1}, ⋯, c_{s}⟩ .

The form ⟨c_{1}, ⋯, c_{s}⟩ is unique up to isometry, and we call it the quasilinear part of φ. The quadratic form φ is called nonsingular if s = 0; singular if s > 0; and totally singular if r = 0.

Let φ be a nonzero quadratic form of dimension n ≥ 1. If φ is not isometric to [0, 0] ⊥ ⟨0, ⋯, 0⟩ and ⟨a, 0, ⋯, 0⟩ for some a ∈ F^{∗}, then the polynomial φ(x_{1}, ⋯, x_{n}) given by φ is irreducible. In this case, we define the function

field of φ, denoted by F (φ), as the quotient field of \frac{F [x_{1}, ⋯, x_{n}]}{(φ(x_{1}, ⋯, x_{n}))}. In other

cases, we set F (φ) = F.

The function field of a bilinear form B is defined as the field F (B̃).

For a field extension K/F and a quadratic form (or a bilinear form) φ, the form φ ⊗ K is denoted by φ_{K}.

Two quadratic forms (or bilinear forms) φ and φ′ are called similar if φ ≃ aφ′ for some scalar a ∈ F^{∗}.

A bilinear form B′ is called a subform of another one B, denoted by B′ ⊂ B, if B ≃ B′ ⊥ B′′ for some bilinear form B′′.

2.1 Witt decomposition of bilinear and quadratic forms

It was proved in [7, Prop. 2.4] that any nonzero quadratic form φ is uniquely decomposed as follows: φ ≃ i×ℍ ⊥ j ×⟨0⟩ ⊥ φ_{an}, where φ_{an} is an anisotropic form called the anisotropic part of φ, and ℍ = [0, 0] is the hyperbolic plane (here n × ψ denotes the orthogonal sum of n copies of a quadratic form ψ). The integer i (resp. j) is called the Witt index of φ and denoted by i_{W} (φ) (resp. the defect index of φ and denoted by i_{d}(φ)). A nonsingular form φ is called hyperbolic if dim φ = 2i_{W} (φ).

Any bilinear form B is decomposed as follows: B ≃ M ⊥ B_{an}, where M is a metabolic bilinear form and B_{an} is an anisotropic bilinear form. The form B_{an} is unique [9], [10], and we call it the anisotropic part of B.

Two quadratic forms (or bilinear forms) φ and φ′ are called Witt-equivalent, denoted by φ ∼ φ′, if φ ⊥ H ≃ φ′ ⊥ H′ where H and H′ are hyperbolic quadratic forms (or metabolic bilinear forms). The condition φ ∼ φ′ implies that φ_{an} ≃ φ′_{an}.

It is clear that a bilinear form B is isotropic if and only if B̃ is isotropic.

2.2 Dominated forms and the subform theorem

Let φ and φ′ be two quadratic forms with underlying vector spaces V and W, respectively. We say that φ is dominated by φ′ (or φ′ dominates φ), denoted by φ ≺ φ′, if there exists an injective F -linear map t : (φ, V) ⟶ (φ′, W) such that φ′(t(v)) = φ(v) for any v ∈ V. We say that φ is weakly dominated by φ′ if aφ ≺ φ′ for some a ∈ F^{∗}.

The following proposition gives an equivalent definition of the domination relation:

Proposition 2.1 ([7, Lem. 3.1]) For quadratic forms φ and φ′, we have an equivalence between:

(1) φ ≺ φ′.

(2) There exist nonsingular forms Q, R, nonnegative integers s′ ≤ s ≤ t, and scalars c_{1}, ⋯, c_{t}, d_{1}, ⋯, d_{s′} ∈ F such that:

(i) φ ≃ R ⊥ ⟨c_{1}, ⋯, c_{s}⟩. (ii) φ′ ≃ Q ⊥ R ⊥ [c_{1}, d_{1}] ⊥ ⋯ ⊥ [c_{s′}, d_{s′}] ⊥ ⟨c_{s′+1}, ⋯, c_{t}⟩ .

The subform theorem asserts the following:

Theorem 2.2 ([7, Th. 4.2]) If φ and φ′ are anisotropic quadratic forms such that φ′ is nonsingular and becomes hyperbolic over F (φ), then φ ≺ aφ′ for any scalar a ∈ D_{F} (φ)D_{F} (φ′).

2.3 On Pfister forms, quasi-Pfister forms and their neighbors

For an n-fold bilinear Pfister form B, the nonsingular form B ⊗ [1, b] is called an (n+1)-fold Pfister form, where ⊗ denotes the action of W (F) on W_{q}(F) [4]. A quasi-Pfister form is a totally singular form associated to a bilinear Pfister form.

We say that a quadratic form φ is a Pfister neighbor (resp. a quasi-Pfister neighbor) if there exists a Pfister form (resp. a quasi-Pfister form) π such that 2 dim φ > dim π and φ is weakly dominated by π. In this case, the form π is unique (up to isometry), and for any field extension K/F we have that φ_{K} is isotropic if and only if π_{K} is isotropic. In particular, π_{F (φ)} and φ_{F (π)} are isotropic [7], [13, Prop. 3.1].

3 Preliminaries on bilinear and quadratic forms

We say that a totally singular form φ is quasi-hyperbolic if dim φ ≥ 2 dim φ_{an}. This definition of quasi-hyperbolicity is different from that fixed in [15], and presents the advantage that it remains invariant under field extensions. However, it should be noted that the results [15, Prop. 1.4, Th. 1.5] that we will use in our proofs, and which are proved using the first definition of quasihyperbolicity, remain true with this new one.

Lemma 3.1 Let B be a bilinear form and φ a quadratic form. We have:

(1) If B is metabolic, then B̃ is quasi-hyperbolic.

(2) If B is anisotropic and B_{F (φ)} is isotropic, then φ is totally singular.

Proof. (1) Let V be the underlying vector space of B. If B is metabolic, then V contains a subspace W of half dimension such that B(W, W) = 0. In particular, B̃(w) = 0 for any w ∈ W, and then dim B̃_{an} ≤ dim V − dim W =

dim W = \frac{dim _{B}}{2}.

(2) Since B̃_{F (φ)} is isotropic and B̃ is anisotropic, the claim follows from [13, Cor. 3.3]. □

Remark 3.2 In general, the quasi-hyperbolicity of the totally singular form B̃ does not imply the metabolicity of B. For example, for x a variable over F and K = F (x), the bilinear form B = ⟨1 : 1 : 0⟩ ⊥ ⟨1 : 0 : x⟩ is not metabolic over K as we can see by using the uniqueness of the anisotropic part, but B̃ ≃ ⟨1, 0, 1, x⟩ ≃ ⟨0, 0, 1, x⟩ is quasi-hyperbolic over K.

Proposition 3.3 An isotropic bilinear Pfister form is metabolic.

Proof. The proposition is obvious for F = ℤ/2ℤ or B of dimension 2. So suppose F ≠ ℤ/2ℤ and dim B > 2. Set B = B′ ⊥ αB′ for some α ∈ F^{∗} and B′ a bilinear Pfister form. If B′ is isotropic, then we conclude by induction on dim B. If not, there exists β ∈ D_{F} (B′) ∩ D_{F} (αB′). By [4, Cor. 2.16, Page 101] B′ is round, i.e. any scalar x ∈ D_{F} (B′) satisfies xB′ ≃ B′. Hence, B′ ≃ βB′ ≃ αB′ and thus B is metabolic. □

We need some results concerning the isotropy of bilinear forms over inseparable quadratic extensions:

Lemma 3.4 Let B be an anisotropic bilinear form and d ∈ F^{∗} − F^{∗2}. We have:

(1) B is isotropic over F (\sqrt{d}) ⟺ there exist a, b ∈ F with a ≠ 0 such that ⟨a : b : ad⟩ ⊂ B.

(2) B is metabolic over F (\sqrt{d}) ⟺ B ≃ ⊥_{i=1}^{r} ⟨a_{i} : b_{i} : a_{i}d⟩ for some a_{i}, b_{i} ∈ F with a_{i} ≠ 0 (1 ≤ i ≤ r).

Proof. Let V be the underlying vector space of B. In both assertions the implication (⟸) is clear.

(1) ⟹: Suppose that B_{F (\sqrt{d})} is isotropic. Then there exist vectors v, v′ ∈ V, not both zero, such that B(v, v) = dB(v′, v′). Since B is anisotropic, we have B(v′, v′) ≠ 0. Moreover, the vectors v and v′ are linearly independent. Otherwise, there would exist α ∈ F^{∗} such that v = αv′. Hence α^{2}B(v′, v′) = dB(v′, v′), and thus d would belong to F^{∗2}. The restriction of B to the space F v ⊕ F v′ is the bilinear form ⟨a : b : ad⟩ where a = B(v′, v′) and b = B(v, v′).

(2) ⟹: Suppose that B_{F (\sqrt{d})} is metabolic. Since B_{F (\sqrt{d})} is isotropic, there exist by statement (1) scalars a_{1}, b_{1} with a_{1} ≠ 0 such that B ≃ ⟨a_{1} : b_{1} : a_{1}d⟩ ⊥ B′ for some bilinear form B′. If dim B = 2, then we are done. If not, we use the uniqueness of the anisotropic part to get that B′_{F (\sqrt{d})} is metabolic. The claim then follows by induction on dim B. □

We get the following corollary:

Corollary 3.5 If B is a bilinear form and d ∈ F^{∗} − F^{∗2}, then there exists a bilinear form B′ defined over F such that (B_{F (\sqrt{d})})_{an} ≃ B′_{F (\sqrt{d})}.

Proof. Without loss of generality, we may suppose that B is anisotropic and becomes isotropic over F (\sqrt{d}). By Lemma 3.4, B ≃⊥_{i=1}^{r} ⟨a_{i} : b_{i} : a_{i}d⟩ ⊥ B′ for some anisotropic bilinear B′ over F (\sqrt{d}), and scalars a_{i}, b_{i} with a_{i} ≠ 0 (1 ≤ i ≤ r). By the uniqueness of the anisotropic part we get (B_{F (\sqrt{d})})_{an} ≃ B′_{F (\sqrt{d})}. □

The following result will be used in the proof of Theorem 1.2:

Proposition 3.6 Let φ and φ′ be totally singular forms such that φ′ is anisotropic of dimension ≥ 2 and dominated by φ. Then we have the inclusion W (F (φ)/F) ⊂ W (F (φ′)/F).

To prove this proposition we need some results from the specialization theory of bilinear and quadratic forms. First, we prove the following lemma:

Lemma 3.7 For φ and φ′ as in Proposition 3.6, there exists an F -place from F (φ) to F (φ′).

Proof. Since φ′ ≺ φ, there exist integers 2 ≤ s ≤ t and scalars c_{1}, ⋯, c_{t} ∈ F such that φ′ ≃ ⟨c_{1}, ⋯, c_{s}⟩ and φ ≃ φ′ ⊥ ⟨c_{s+1}, ⋯, c_{t}⟩. Without loss of generality, we may suppose c_{1} = 1. Since dim φ_{an} ≥ dim φ′ ≥ 2, the homogeneous polynomials given by φ and φ′ are irreducible and we have:

F (φ) = F (x_{2}, ⋯, x_{t})(\sqrt{α})
F (φ′) = F (y_{2}, ⋯, y_{s})(\sqrt{α′}),

where α = ∑_{i=2}^{t} c_{i}x_{i}^{2} and α′ = ∑_{i=2}^{s} c_{i}y_{i}^{2}. By [19, Cor. 6.13, page 162], there exists an F -place λ from F (x_{2}, ⋯, x_{t}) to F (y_{2}, ⋯, y_{s}) given by:

λ(x_{i}) = \begin{cases}y_{i} & if 2 ≤ i ≤ s \\ 0 & otherwise. \end{cases}

Since λ(α) = α′, the F -place λ restricts to an F^{2}-place from F^{2}(x_{2}^{2}, ⋯, x_{t}^{2})(α) to F^{2}(y_{2}^{2}, ⋯, y_{s}^{2})(α′). Moreover, the squaring map (resp. its inverse) yields field isomorphisms F (φ) ⟶ F^{2}(x_{2}^{2}, ⋯, x_{t}^{2})(α) and F^{2}(y_{2}^{2}, ⋯, y_{s}^{2})(α′) ⟶ F (φ′). It is now clear that composing these isomorphisms with λ yields the desired F -place from F (φ) to F (φ′). □

We also need a specialization result by Knebusch:

Proposition 3.8 ([11, Th. 3.1]) Let K and L be fields (of any characteristic), and let λ : K ⟶ L∪{∞} be a place. Then there exists a unique additive map λ_{∗} : W (K) ⟶ W (L) defined as follows: λ_{∗}(⟨a⟩_{b}) = ⟨λ(a)⟩_{b} for every a ∈ K such that λ(a) ≠ 0, ∞, and λ_{∗}(⟨a⟩_{b}) = 0 for every a ∈ K such that λ(ac^{2}) = 0 or ∞ for every c ∈ K^{∗}.

Now we are able to prove Proposition 3.6:

Proof. By Lemma 3.7 there exists an F -place λ : F (φ) ⟶ F (φ′) ∪ {∞}. This place induces an additive map λ_{∗} : W (F (φ)) ⟶ W (F (φ′)) defined as in Proposition 3.8. Since λ_{∗}(⟨a⟩_{b}) = ⟨λ(a)⟩_{b} = ⟨a⟩_{b} for every a ∈ F^{∗}, the claim follows. □

We give an analogue of Proposition 3.6 for Witt kernels for quadratic forms:

Proposition 3.9 Let φ and φ′ be quadratic forms such that φ is anisotropic and becomes isotropic over F (φ′). Then W_{q}(F (φ)/F) ⊂ W_{q}(F (φ′)/F).

Proof. Suppose W_{q}(F (φ)/F) ≠ 0, and let η ∈ W_{q}(F (φ)/F) be anisotropic. We proceed by induction on dim η. By the subform theorem φ is weakly dominated by η, and thus η_{F (φ′)} is isotropic since φ_{F (φ′)} is also isotropic. Hence F (φ′)(η)/F (φ′) is purely transcendental.

(1) If η is similar to a Pfister form, then η ∈ W_{q}(F (φ′)/F) and we are done.

(2) If η is not similar to a Pfister form, then η_{1} := (η_{F (η)})_{an} ≠ 0 [14]. The form φ_{F (η)} is anisotropic, otherwise F (η)(φ)/F (η) would be purely transcendental [13, Cor. 3.4], and thus η_{F (η)} would be hyperbolic since η_{F (φ)} ∼ 0. Since φ_{F (η)(φ′)} is isotropic and (η_{1})_{F (η)(φ)} ∼ η_{F (η)(φ)} ∼ 0, we deduce by induction that (η_{1})_{F (η)(φ′)} ∼ 0. Since F (η)(φ′) = F (φ′)(η) and F (φ′)(η)/F (φ′) is purely transcendental, the form η_{F (φ′)} is then hyperbolic. Hence the claim. □

Remark 3.10 Note that in this proof, the use of the subform theorem is necessary when φ′ is totally singular. However, for φ′ not totally singular, the subform theorem can be avoided since in this case the isotropy of φ_{F (φ′)} implies that F (φ′)(φ)/F (φ′) is purely transcendental [13, Cor. 3.4], and thus W_{q}(F (φ)/F) ⊂ W_{q}(F (φ′)/F).

4 Proofs

4.1 Proof of Proposition 1.1

Let B be an anisotropic bilinear form and φ an anisotropic quadratic form such that B is metabolic over F (φ). Let α = uv with u ∈ D_{F} (φ) and v ∈ D_{F} (B) = D_{F} (B̃). Since B_{F (φ)} is metabolic, B̃_{F (φ)} is quasi-hyperbolic and φ is totally singular (Lemma 3.1), hence it follows from [15, Prop. 1.4] that uφ ⊂ vB̃, hence D_{F} (φ) ⊂ αD_{F} (B̃) = αD_{F} (B). Write φ = ⟨a_{1}, ⋯, a_{n}⟩, and let x_{i} be elements in the underlying vector space V of B such that B(x_{i}, x_{i}) = α^{−1}a_{i}. The x_{i} are F -linearly independent since the a_{i} are F^{2}-linearly independent because of the anisotropy of φ. Let W be the subspace of V generated by the x_{i}, and B′ ≃ αB |_{W}. By the anisotropy of B, we conclude that there exists a bilinear form B′′ with αB ≃ B′ ⊥ B′′. Clearly, D_{F} (B′) = D_{F} (φ), dim B′ = dim φ, implying the result. □

4.2 Proof of Theorem 1.2

We have to show that if φ is an anisotropic quadratic form of dimension ≥ 2 such that W (F (φ)/F) ≠ 0, then φ is totally singular, and any anisotropic B in W (F (φ)/F) ≠ 0 decomposes into an orthogonal sum of forms similar to bilinear Pfister forms in A(φ_{qp}).

The metabolicity of B_{F (φ)} implies that φ is totally singular. We may suppose that φ represents 1, and that ndeg_{F} (φ) = 2^{d}. We will use induction on d. Suppose d = 1. Then necessarily φ ≃ ⟨1, a⟩ where N_{F} (φ) = F^{2}(a). In this situation, F (φ) = F (\sqrt{a})(t) for some transcendental element t, and thus B becomes already metabolic over F (\sqrt{a}). The theorem now follows immediately from Lemma 3.4(2).

So suppose d ≥ 2. Then we may write φ = ⟨1, a_{1}, ⋯, a_{d}, ⋯ ⟩ such that N_{F} (φ) = F^{2}(a_{1}, ⋯, a_{d}). Let φ′ = ⟨1, a_{1}, ⋯, a_{d−1}⟩. Then N_{F} (φ′) = F^{2}(a_{1}, ⋯, a_{d−1}), ndeg_{F} (φ′) = 2^{d−1}. By Proposition 3.6, B_{F (φ′)} is metabolic. By induction, B ≃⊥_{i=1}^{m} x_{i}C_{i} with (d − 1)-fold bilinear Pfister forms C_{i} associated to φ′_{qp}. In particular, D_{F} (C_{i}) ∪ {0} = D_{F} (φ′_{qp}) ∪ {0} = N_{F} (φ′) = F^{2}(a_{1}, ⋯, a_{d−1}).

After scaling, we may suppose that x_{1} = 1. If we write C′ = C_{1} = ⟨1, c_{1}⟩_{b}⊗⋯⊗ ⟨1, c_{d−1}⟩_{b}, then D_{F} (C′)∪{0} = F^{2}(c_{1}, ⋯, c_{d−1}) = F^{2}(a_{1}, ⋯, a_{d−1}) = N_{F} (φ′), and we also can write B ≃ C′ ⊥ B′ for some bilinear form B′.

Since B is metabolic over F (φ), we have that B̃ is quasi-hyperbolic over F (φ).

By [15, Prop. 1.4] and since B represents 1, we have N_{F} (φ) ⊂ D_{F} (B) ∪ {0} = D_{F} (B̃) ∪ {0}, and by [15, Th. 1.5], we know that dim B = dim B̃ is a multiple of 2^{d}.

So D_{F} (C′) ∪ {0} = N_{F} (φ′) ⊂ N_{F} (φ) ⊂ D_{F} (C′ ⊥ B′) ∪ {0}, and since ndeg_{F} (φ′) = 2^{d−1} < 2^{d} = ndeg_{F} (φ), there exists u ∈ D_{F} (C′) ∪ {0} and c_{d} ∈ D_{F} (B′)∪{0} such that N_{F} (φ′)(u+c_{d}) = F^{2}(c_{1}, ⋯, c_{d−1}, u+c_{d}) = N_{F} (φ). Necessarily, c_{d} ≠ 0 since u ∈ N_{F} (φ′), and we have N_{F} (φ′)(u + c_{d}) = N_{F} (φ′)(c_{d}) = N_{F} (φ).

Now let C = C′ ⊗ ⟨1, c_{d}⟩_{b}. Note that D_{F} (C) ∪ {0} = N_{F} (φ) = N_{F} (φ_{qp}) = D_{F} (φ_{qp}) ∪ {0} and dim C = dim φ_{qp} = 2^{d}, implying that C is associated to φ_{qp}.

Consider C ⊥ B. Since C′ ⊥ ⟨c_{d}⟩_{b} ⊂ B and C′ ⊥ ⟨c_{d}⟩_{b} ⊂ C, we have that C ⊥ B ≃ M ⊥ B′′ with B′′ anisotropic and M metabolic of dimension ≥ 2^{d} + 2 = 2 dim(C′ ⊥ ⟨c_{d}⟩_{b}). By dimension count, dim B′′ < dim B. Now C_{F (φ)} is clearly isotropic and hence metabolic, so that B′′_{F (φ)} is metabolic as well, and therefore its dimension is a multiple of 2^{d} [15, Th. 1.5]. In particular, dim B′′ ≤ dim B − 2^{d} as dim B is a multiple of 2^{d}. In W (F), we therefore have B ∼ C ⊥ C ⊥ B ∼ C ⊥ B′′ with B anisotropic and dim(C ⊥ B′′) ≤ dim B. Consequently, B ≃ C ⊥ B′′, and the theorem follows by an easy induction on the dimension of B. □

4.3 Proof of Theorem 1.4

Let φ be an anisotropic Pfister neighbor or quasi-Pfister neighbor of π.

(1) Since the forms φ_{F (π)} and π_{F (φ)} are isotropic, it follows from Proposition 3.9 that W_{q}(F (φ)/F) = W_{q}(F (π)/F). Hence, for the proofs of statements (2)-(3) we suppose that φ ≃ π.

(2) Suppose that φ is a Pfister form. Let ψ ∈ W_{q}(F (φ)/F) be anisotropic. By the subform theorem x_{1}φ ≺ ψ for some x_{1} ∈ F^{∗}. If dim ψ = dim φ, we take τ = ⟨x_{1}⟩_{b} and we are done. If not, we write ψ ≃ x_{1}φ ⊥ ψ′ for some nonsingular form ψ′. Since ψ_{F (φ)} is hyperbolic, the form ψ′_{F (φ)} is also hyperbolic. Since dim ψ′ < dim ψ, we get by induction on dim ψ that ψ′ ≃ x_{2}φ ⊥ ⋯ ⊥ x_{r}φ for some x_{2}, ⋯, x_{r} ∈ F^{∗}. Hence, ψ ≃ τ ⊗ φ where τ = ⟨x_{1}, ⋯, x_{r}⟩_{b}.

(3) Suppose that φ is a quasi-Pfister form. Let ψ ∈ W_{q}(F (φ)/F) be anisotropic, and let B = ⟨1, a_{1}⟩_{b} ⊗ ⋯ ⊗ ⟨1, a_{n}⟩_{b} be such that φ ≃ B̃. By [8, Rem. 5.2(i)], ψ ∼ B ⊗ ρ ∈ W_{q}(F) for some nonsingular form ρ which we may suppose of minimal dimension among all γ with ψ ∼ B ⊗γ. In particular, ρ is anisotropic. Let x_{i} ∈ F^{∗} and b_{i} ∈ F such that ρ ≃⊥_{i=1}^{r} x_{i}[1, b_{i}]. Suppose that B ⊗ ρ is isotropic. Then, after reindexing, there exist 2 ≤ s ≤ r, y_{i} ∈ D_{F} (x_{i}B ⊗ [1, b_{i}]), 1 ≤ i ≤ s, such that y_{1} + ⋯ + y_{s} = 0. By roundness of the Pfister forms B ⊗ [1, b_{i}], x_{i}B ⊗ [1, b_{i}] ≃ y_{i}B ⊗ [1, b_{i}] for i ≤ s. Let y_{i} := x_{i} for i > s and put ρ′ =⊥_{i=1}^{r} y_{i}[1, b_{i}]. By the above, ψ ∼ B ⊗ ρ ≃ B ⊗ ρ′, but ρ′ is obviously isotropic, a contradiction. Hence B ⊗ ρ is anisotropic and thus ψ ≃ B ⊗ ρ. □

4.4 Proof of Theorem 1.5

Let φ be an anisotropic quadratic form of dimension ≥ 3, and let φ′ be a quadratic form dominated by φ such that dim φ = dim φ′+1 and W_{q}(F (φ′)/F) is a strong n-Pfister group for some n ≥ 1. We have to prove that W_{q}(F (φ)/F) is an {n, n + 1}-Pfister group.

Let ψ ∈ W_{q}(F (φ)/F). We use induction on dim ψ, the initial step dim ψ ≤ 2 being trivial. So we may assume ψ anisotropic of dimension > 2. Since ψ ∈ W_{q}(F (φ′)/F) (Proposition 3.9) and by assumption, there exist forms ρ_{i} ∈ GP_{n}F ∩ W_{q}(F (φ′)/F) such that ψ ≃⊥_{i=1}^{r} ρ_{i}. After scaling, we may assume that φ′ and ρ_{1} represent 1, so that φ′ ≺ ρ_{1}. Put γ =⊥_{i=2}^{r} ρ_{i}.

Now ρ_{1} is hyperbolic over F (φ) if and only if φ is weakly dominated by ρ_{1} if and only if ρ_{1} ∈ GP_{n}F ∩ W_{q}(F (φ)/F), in which case γ_{F (φ)} is hyperbolic, and we are done by induction. So we may assume that ρ_{1} does not dominate a form similar to φ.

Since ρ_{1} dominates φ′ but not φ, it follows that i_{t}(φ ⊥ ρ_{1}) = dim φ′ (cf. [7, Cor. 3.13]; here i_{t} denotes the total index i_{W} +i_{d}). If σ denotes the anisotropic part of φ ⊥ ρ_{1}, then φ ⊥ ρ_{1} ≃ u × ⟨0⟩ ⊥ v × ℍ ⊥ σ with u + v = dim φ′.

On the other hand, since ψ is hyperbolic over F (φ), and since both ψ and φ represent 1, we have that ψ dominates φ, hence i_{t}(φ ⊥ ψ) = dim φ = dim φ′ + 1. Therefore,

φ ⊥ ψ ≃ u × ⟨0⟩ ⊥ v × ℍ ⊥ σ ⊥ γ,

and comparing the total index on both sides shows that σ ⊥ γ is isotropic, so that there exists x ∈ D_{F} (σ) ∩ D_{F} (γ) since σ and γ are anisotropic.

Now consider π = ρ_{1} ⊥ xρ_{1} which is anisotropic because it dominates ρ_{1} ⊥ ⟨x⟩ which is a Pfister neighbor of π and in turn dominated by the anisotropic form ψ. Then φ ⊥ π ≃ u × ⟨0⟩ ⊥ v × ℍ ⊥ σ ⊥ xρ_{1} with σ ⊥ xρ_{1} isotropic. Hence, i_{t}(φ ⊥ π) ≥ dim φ′ + 1 = dim φ, which by [7, Cor. 3.13] implies equality and that π dominates φ. In particular, π ∈ P_{n+1}F ∩ W_{q}(F (φ)/F).

Now let ψ′ be the anisotropic part of ψ ⊥ π. We have that both ψ and π dominate ρ_{1} ⊥ ⟨x⟩, so i_{t}(ψ ⊥ π) ≥ dim(ρ_{1} ⊥ ⟨x⟩) = 2^{n} + 1, and since i_{t}(ψ ⊥ π) = i_{W} (ψ ⊥ π) by nonsingularity, we have that dim ψ′ ≤ dim ψ + dim π − 2(2^{n} + 1) < dim ψ. But clearly, ψ′ ∈ W_{q}(F (φ)/F), and the proof readily concludes by induction on the dimension. □

4.5 Proof of Theorem 1.6

Let φ be an anisotropic quadratic form of dimension ≥ 3, and let φ′ be a Pfister neighbor of an n-fold Pfister form such that φ′ ≺ φ and dim φ = dim φ′ + 1.

If φ is a Pfister neighbor of an n-fold Pfister form, then W_{q}(F (φ)/F) is a strong n-Pfister group (Theorem 1.4(2)). So suppose that φ is not a Pfister neighbor of an n-fold Pfister form. Our aim is to prove that W_{q}(F (φ)/F) is a strong (n + 1)-Pfister group. Let ρ be the Pfister form associated to φ′ and let ψ ∈ W_{q}(F (φ)/F) be anisotropic. Then ψ becomes hyperbolic over F (φ′) and thus over F (ρ). After scaling, we may assume that ψ and φ′ represent 1, so that by Theorem 1.4(2), ψ ≃ ρ ⊥ τ ⊗ ρ for some bilinear form τ. We now repeat the same proof as for Theorem 1.5, with ρ_{1} replaced by ρ and γ by τ ⊗ ρ, by noting that ρ does not dominate a form similar to φ because by assumption φ is not a Pfister neighbor of some n-fold Pfister form. We also let σ be the anisotropic part of φ ⊥ ρ.

As before, it follows that there exists x ∈ D_{F} (σ) ∩ D_{F} (τ ⊗ ρ). By a standard argument, the roundness of ρ readily implies that τ ⊗ ρ ≃ xρ ⊥ τ′ ⊗ ρ for some bilinear form τ′.

Now consider π = ρ ⊥ xρ ∈ P_{n+1}F. As above, we have that φ is dominated by π. Hence, π ∈ P_{n+1}F ∩ W_{q}(F (φ)/F). Therefore also τ′ ⊗ ρ ∈ W_{q}(F (φ)/F), and the proof can now readily be completed by induction on dim ψ as ψ ≃ π ⊥ τ′ ⊗ ρ. □

5 Pfister neighbors for bilinear forms

In view of the subform theorem given in Proposition 1.1, we suggest the following definition for Pfister neighbors in the case of bilinear forms:

Definition 5.1 We say that a bilinear form B′ is a Pfister neighbor of a bilinear Pfister form B if 2 dim B′ > dim B and there exists B′′ ∈ A(B̃′) similar to a subform of B.

As for quadratic forms we prove the following:

Proposition 5.2 Let B′ be a Pfister neighbor of a bilinear Pfister form B.

Then:

(1) For any field extension K/F, the forms B_{K} and B′_{K} are simultaneously isotropic or anisotropic. In particular, B_{F (B′)} and B′_{F (B)} are isotropic.

(2) If B′ is anisotropic and a Pfister neighbor of another bilinear Pfister form ρ, then B̃ ≃ ρ̃.

Proof. (1) Let B′′ ∈ A(B̃′) similar to a subform of B. The forms B′_{K} and B′′_{K} are simultaneously isotropic or anisotropic since B̃′′ ≃ B̃′. If B_{K} is isotropic, then it is metabolic and B′′_{K} becomes isotropic as dim B′′ = dim B′ > \frac{dim B}{2} . Obviously, the isotropy of B′′_{K} implies that of B_{K}.

(2) By statement (1) B′_{F (ρ)} is isotropic and thus B_{F (ρ)} is also isotropic. Since B is anisotropic, dim B = dim ρ and 1 ∈ D_{F} (B) ∩ D_{F} (ρ), it follows from Proposition 1.1 that B ∈ A(ρ̃), i.e. B̃ ≃ ρ̃. □

We give an equivalent definition to 5.1:

Proposition 5.3 Let B and B′ be anisotropic bilinear forms such that B is a bilinear Pfister form. Then B′ is a Pfister neighbor of B if and only if B becomes isotropic over F (B′) and 2 dim B′ > dim B.

Proof. If B is isotropic over F (B′), then it is metabolic over F (B′), and by Proposition 1.1 there exists a bilinear form B′′ ∈ A(B̃′) similar to a subform of B. If moreover 2 dim B′ > dim B, then B′ is a Pfister neighbor of B. The converse follows from Proposition 5.2(1). □

Theorem 1.2 allows us important corollaries:

Corollary 5.4 Let B be an anisotropic bilinear form and φ an anisotropic totally singular form. Suppose that B_{F (φ)} is metabolic and 2 dim φ > dim B. Then B is similar to a bilinear Pfister form ρ, and any bilinear form B′ ∈ A(φ) is a Pfister neighbor of ρ.

Proof. By Theorem 1.2, B ≃⊥_{i=1}^{r} α_{i}B_{i} for some scalars α_{i} ∈ F^{∗} and dfold bilinear Pfister forms B_{i} associated to φ_{qp}. Since dim φ_{qp} ≥ dim φ and 2 dim φ > dim B = r dim φ_{qp}, it follows that r = 1 and thus B is similar to the bilinear Pfister form ρ = B_{1}. Moreover, it follows from Proposition 5.3 that any B′ ∈ A(φ) is a Pfister neighbor of ρ since 2 dim B′ = 2 dim φ > dim B = dim ρ and ρ is isotropic over F (φ) = F (B′). □

This corollary allows us to classify anisotropic bilinear forms of height 1. This makes up the first step for a standard splitting theory of bilinear forms:

Corollary 5.5 An anisotropic bilinear form B becomes metabolic over F (B) if and only if it is similar to a bilinear Pfister form.

Proof. We use Proposition 3.3, and Corollary 5.4 applied to the case φ = B̃. □

Corollary 5.6 Let B be an anisotropic bilinear form. Then the following are equivalent:

(1) B is a Pfister neighbor of some bilinear Pfister form.

(2) There exist anisotropic bilinear forms B′, B′′ such that dim B′′ < dim B′, B′ ∈ A(B̃) and B′_{F (B)} ∼ B′′_{F (B)}.

(3) There exists a bilinear form B′ ∈ A(B̃) such that the anisotropic part of B′_{F (B)} is defined over F.

Proof. (2) ⟹ (1) The assumption B′_{F (B)} ∼ B′′_{F (B)} implies (B′ ⊥ B′′)_{F (B)} ∼ 0. Since B′ is anisotropic and dim B′′ < dim B′, the bilinear form B′ ⊥ B′′ can not be metabolic. Moreover, 2 dim B̃ = 2 dim B′ > dim(B′ ⊥ B′′) ≥ dim(B′ ⊥ B′′)_{an}. We apply Corollary 5.4 with φ = B̃ to get that (B′ ⊥ B′′)_{an} is similar to a bilinear Pfister form ρ, and B is a Pfister neighbor of ρ since B ∈ A(B̃).

(3) ⟹ (2) If B′′ is a bilinear form over F satisfying (B′_{F (B)})_{an} ≃ B′′_{F (B)}, then clearly B′_{F (B)} ∼ B′′_{F (B)}, and dim B′′ < dim B′ since B′_{F (B)} is isotropic.

(1) ⟹ (3) Let B′ and B′′ be bilinear forms such that B′ ∈ A(B̃), B′ ⊥ B′′ is similar to a d-fold bilinear Pfister form and dim B > 2^{d−1}. Since B′_{F (B)} is isotropic, we get B′_{F (B)} ∼ B′′_{F (B)}. Moreover, dim B′′ = 2^{d} − dim B′ = 2^{d} − dim B < 2^{d−1} < dim B. By [8, Th. 1.1], B′′_{F (B)} is anisotropic and thus (B′_{F (B)})_{an} ≃ B′′_{F (B)}. □

6 Other results on Witt kernels

We begin this section by a criterion about the metabolicity over inseparable biquadratic extensions. The case of quadratic extensions is covered by Theorem 1.2.

Proposition 6.1 Let B be an anisotropic bilinear form of dimension ≥ 2 and a_{1}, a_{2} ∈ F^{∗}. Then the following are equivalent:

(1) B is metabolic over F (\sqrt{a_{1}}, \sqrt{a_{2}}).

(2) B ∼ (⊥_{i=1}^{r} α_{i}λ_{i}) ⊥ (⊥_{i=1}^{s} β_{i}μ_{i}) for some scalars α_{i}, β_{j} ∈ F^{∗}, 1-fold bilinear Pfister forms λ_{i} ∈ A(⟨1, a_{1}⟩), and μ_{j} = ⟨1, x_{j} + a_{2}⟩_{b} with x_{j} ∈ F^{2}(a_{1}).

Proof. (1) ⟹ (2) By Theorem 1.2, it suffices to prove the existence of scalars β_{j} ∈ F^{∗}, and bilinear forms μ_{j} = ⟨1, x_{j} + a_{2}⟩_{b} with x_{j} ∈ F^{2}(a_{1}) such that (B ⊥ (⊥_{j=1}^{s} β_{j}μ_{j}))_{F (\sqrt{a_{1}})} ∼ 0. We proceed by induction on dim B and we may suppose that B_{F (\sqrt{a_{1}})} is not metabolic.

By Corollary 3.5 there exists an anisotropic bilinear form B′ over F such that (B_{F (\sqrt{a_{1}})})_{an} ≃ B′_{F (\sqrt{a_{1}})}. Let β_{1} ∈ D_{F} (B′). Since B′_{F (\sqrt{a_{1}})} is metabolic over F (\sqrt{a_{1}}, \sqrt{a_{2}}), it follows from Proposition 1.1 that B′_{F (\sqrt{a_{1}})} ≃ β_{1}B′_{1} ⊥ B′_{2} for some B′_{1} ∈ A(⟨1, a_{2}⟩_{F (\sqrt{a_{1}})}) and a bilinear form B′_{2} over F (\sqrt{a_{1}}). Let u, v be vectors in the underlying vector space V of B′_{1} such that B′_{1}(u, u) = 1 and B′_{1}(v, v) = a_{2}. Since ⟨1, a_{2}⟩ is anisotropic, the vectors u and v are F -linearly independent. Let α = B′_{1}(u, v). By considering the basis {u, αu + v} of V, we get the isometry B′_{1} ≃ ⟨1, x_{1} + a_{2}⟩_{b}, where x_{1} = α^{2} ∈ F^{2}(a_{1}). In particular, (B′ ⊥ β_{1} ⟨1, x_{1} + a_{2}⟩_{b})_{F (\sqrt{a_{1}})} ∼ B′_{2}.

If dim B = 2, then B′_{2} = 0 since we may take B = B′. So suppose dim B > 2 and B′_{2} ≁ 0. Since ((B′ ⊥ β_{1} ⟨1, x_{1} + a_{2}⟩_{b})_{F (\sqrt{a_{1}})})_{an} ≃ B′_{2}, there exists by Corollary 3.5 an anisotropic bilinear form B′′ over F such that B′_{2} ≃ B′′_{F (\sqrt{a_{1}})}. The form B′′_{F (\sqrt{a_{1}})} is metabolic over F (\sqrt{a_{1}}, \sqrt{a_{2}}) since B′ is also metabolic over F (\sqrt{a_{1}}, \sqrt{a_{2}}) and x_{1} + a_{2} is a square in F (\sqrt{a_{1}}, \sqrt{a_{2}}). Since dim B′′ < dim B, we conclude by induction.

(2) ⟹ (1) The forms λ_{i} are metabolic over F (\sqrt{a_{1}}), and for any j ∈ \{1, ⋯, s\} the form μ_{j} is also metabolic over F (\sqrt{a_{1}}, \sqrt{a_{2}}) since x_{j} + a_{2} is a square in this field. □

Theorem 1.6 generalizes [6, Prop. 1.4], and together with Theorem 1.5 it implies the following corollary which is a partial generalization of [6, Th. 2.1]:

Corollary 6.2 Let φ′ be a codimension 1 Pfister neighbor of an n-fold Pfister form, and let φ be a quadratic form which is not a Pfister neighbor and dominates φ′ such that dim φ = dim φ′ +2. Then W_{q}(F (φ)/F) is an (n+2)-Pfister group.

Proof. We reproduce the same proof like that for [6, Th. 2.1]. We have W_{q}(F (φ)/F) ∩ P_{n+1}F = 0 since dim φ = 2^{n} + 1 and φ is not a Pfister neighbor. Let φ′′ be a quadratic form such that φ′ ≺ φ′′ ≺ φ and dim φ′′ = dim φ′ + 1. The form φ′′ is not a Pfister neighbor, otherwise φ′′ ∈ GP_{n}F since dim φ′′ = dim φ′ + 1 = 2^{n}, and thus φ would be a Pfister neighbor. By Theorem 1.6, W_{q}(F (φ′′)/F) is a strong (n + 1)-Pfister group, and by Theorem 1.5 W_{q}(F (φ)/F) is an (n + 2)-Pfister group since W_{q}(F (φ)/F) ∩ P_{n+1}F = 0. □

We combine [1, Cor. 2.8], Theorems 1.4, 1.5, 1.6 and Corollary 6.2 to describe W_{q}(F (φ)/F) for φ of small dimension:

Corollary 6.3 Let φ be an anisotropic quadratic form of dimension at most 5, and let s denote the dimension of its quasilinear part. Then we have the following description:

Dimension of φ  Condition on φ  The kernel W_{q}(F (φ)/F)
2  s = 0  strong 1-Pfister group
s = 2  strong 2-Pfister group
3 s = 1 strong 2-Pfister group
s = 3  strong 3-Pfister group
s = 0 and φ is not similar  strong 3-Pfister group

to a 2-fold Pfister form

s = 2  strong 3-Pfister group
4 s = 4 and φ is similar to strong 3-Pfister group

a 2-fold quasi-Pfister form

s = 0 and φ is similar  strong 2-Pfister group

to a 2-fold Pfister form

s = 4 and φ is not similar  {3, 4}-Pfister group

to a 2-fold quasi-Pfister form

s = 1 or 3 and φ is not  4-Pfister group

a Pfister neighbor

5 s = 1 or 3 and φ is strong 3-Pfister group

a Pfister neighbor

s = 5 and φ is  strong 4-Pfister group

quasi-Pfister neighbor

s = 5 and φ is not a  ?

quasi-Pfister neighbor

Proof. Set φ ≃ [a_{1}, b_{1}] ⊥ ⋯ ⊥ [a_{r}, b_{r}] ⊥ ⟨c_{1}, ⋯, c_{s}⟩. By [1, Cor. 2.8], we may exclude the case dim φ = s = 2, and by Theorem 1.4 we may exclude the case where φ is a Pfister neighbor or a quasi-Pfister neighbor. It remains to treat the following cases:

(1) dim φ = 4 and r > 0. We apply Theorem 1.6 to φ′ = [a_{1}, b_{1}] ⊥ ⟨a_{2}⟩ or [a_{1}, b_{1}] ⊥ ⟨c_{1}⟩ according as s = 0 or s = 2.

(2) dim φ = 4 and r = 0. We apply Theorem 1.5 to φ′ = ⟨c_{1}, c_{2}, c_{3}⟩ since we know that W_{q}(F (φ′)/F) is a strong 3-Pfister group.

(3) dim φ = 5 and r = 1 or 2. We apply Corollary 6.2 to φ′ = [a_{1}, b_{1}] ⊥ ⟨c_{1}⟩. □

References

[1] H. Ahmad, On quadratic forms isotropic over inseparable quadratic extensions, Arch. Math. (Basel) 63 (1994) 23–29.

[2] H. Ahmad, Witt kernels of bi-quadratic extensions in characteristic 2, preprint.

[3] R. Baeza, Ein Teilformensatz für quadratische Formen in Charakteristik 2, Math. Z. 135 (1974) 175–184.

[4] R. Baeza, Quadratic forms over semilocal rings, Lect. Notes Math. vol. 655, Berlin, Heidelberg, New York: Springer 1978.

[5] R. Elamn, T. Y. Lam, A. Wadsworth, Pfister Ideals in Witt Rings, Math. Ann. 245 (1979) 219–245.

[6] R. W. Fitzgerald, Witt kernels of function field extensions, Pacific J. Math. 109 (1983) 89–106.

[7] D. W. Hoffmann, A. Laghribi, Quadratic forms and Pfister neighbors in characteristic 2, Trans. Amer. Math. Soc. 356 (2004) 4019-4053.

[8] D. W. Hoffmann, A. Laghribi, Isotropy of quadratic forms over the function field of a quadric in characteristic 2, to appear in J. Algebra.

[9] D. Husemoller, J. Milnor, Symmetric bilinear forms, New York, Heidelberg: Springer 1973.

[10] M. Knebusch, Grothendieck und Wittringe von nichtausgearteten symmetrischen Bilinearformen, S.-B. Heidelberger Akad. Wiss. Math.-Natur. Kl. (1969-1970) 93–157.

[11] M. Knebusch, Specialization of quadratic and symmetric bilinear forms, and a norm theorem, Acta Arithm. 24 (1973) 279–299.

[12] M. Knebusch, Generic splitting of quadratic forms II, Proc. London Math. Soc. 34 (1977) 1–31.

[13] A. Laghribi, Certaines formes quadratiques de dimension au plus 6 et corps des fonctions d’une quadrique en caractéristique 2, Israel J. Math. 129 (2002) 317-361.

[14] A. Laghribi, On the generic splitting of quadratic forms in characteristic 2, Math. Z. 240 (2002) 711-730.

[15] A. Laghribi, Quasi-hyperbolicity of totally singular quadratic forms, Contemp. Math. 344 (2004) 237-248.

[16] A. Laghribi, Witt kernels of quadratic forms for purely inseparable multiquadratic extensions in characteristic 2, Preprint 2004.

[17] P. Mammone, R. Moresi, Formes quadratiques, algèbres à division et extensions multiquadratiques inséparables, Bull. Belg. Math. Soc. 2 (1995) 311–319.

[18] C.-H. Sah, Symmetric bilinear forms and quadratic forms, J. Algebra 20 (1972) 144–160.

[19] W. Scharlau, Quadratic and Hemitian forms, Grundlehren Math. Wiss. Bd. 270, Berlin, Heidelberg, New York, Tokyo: Springer 1985.